A ball is thrown vertically upward from the top of a cliff. The height of the ball is modeled by the function h(t) = 65 + 10t - 5t^2, where h(t) is the height in metres and t is the time in seconds. Determine when the ball reaches its maximum height.

Answer by nerdybill(7384)

A ball is thrown vertically upward from the top of a cliff. The height of the ball is modeled by the function h(t) = 65 + 10t - 5t^2, where h(t) is the height in metres and t is the time in seconds. Determine when the ball reaches its maximum height.
.
equation is
h(t) = 65 + 10t - 5t^2
h(t) = -5t^2 + 10t + 65
the max height is reached at the vertex.
the value of t at the vertex is:
t = -b/(2a)
t = -10/(2(-10))
t = -10/(-20)
t = 10/(20)
t = 1/2 seconds (answer)